a projctile is fired at the angle of 53 degree with the horizontal. the speed of the projectile is 200m/s calculate range (use cos53=0.8,sin53=0.6)
1
Expert's answer
2021-03-24T19:49:38-0400

Let's first find the time that the projectile takes to reach its maximum height:


"v_y=v_0sin\theta-gt,""0=v_0sin\theta-gt""t=\dfrac{v_0sin\theta}{g}."

Then, we can find the total flight time of the projectile:


"t_{flight}=2t=\dfrac{2v_0sin\theta}{g}."

Finally, we can find the range of the projectile:


"x=v_0tcos\theta,""x=v_0\dfrac{2v_0sin\theta}{g}cos\theta=\dfrac{v_0^2sin2\theta}{g},""x=\dfrac{(200\ \dfrac{m}{s})^2sin2\cdot53^{\circ}}{9.8\ \dfrac{m}{s^2}}=3923.5\ m."